OpenStudy (anonymous):

2x squared + 3x + 1 Into equivilant factored form

OpenStudy (anonymous):

1. 2x^2+3x+1 2x^2 + 2x + x + 1 2x(x+1) + (x+1) (x+1)(2x+1

OpenStudy (anonymous):

\[2 x^2+3x+1\] \[2x^2+2x+x+1\] \[2x \left( x+1 \right)+1\left( x+1 \right)\] \[\left( x+1 \right)\left( 2x+1 \right)\]